OpenStudy (anonymous):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bars sell for $5 per pound and the milk chocolate fudge bars sell for $4 per pound. You purchased a combination of these bars for your family and friends and your bill came to $150. This scenario can be modeled by the equation 5d + 4m = 150. Which of the following is not a possible combination of dark chocolate fudge bars and milk chocolate fudge bars? Answer 12 dark bars and 23 milk bars 14 dark bars and 20 milk bars 6 dark bars and 30 milk bars

OpenStudy (anonymous):

You just plug in the values in each solution set to see if it equals 150. So for the first solution, plug in 12 for d and 23 for m. So you now have (12)(5)+(23)(4)=150 This reduces to 60+92=150 Since 152 does not equal 150, it isn't a solution. Check the other possible solutions You have (14)(5)+(20)(4)=150 This reduces to 70+80=150 Since 150=150 this is a solution. Checking the last possible solution. You have (6)(5)+(30)(4)=150 This reduces to 30+120=150 Since 150=150 this is a solution, also.